OpenStudy (anonymous):

Jennifer makes large cookies. These cookies are 14 inches in diameter. How much icing will she need to cover the top of each cookie?

OpenStudy (anonymous):

14π in2 49π in2 196π in2 7π in2

OpenStudy (jewlzme17):

----> Answer: 49π in2 Icing on cookies indicates that this is an area of a circle problem. A = pi*r^2 A = pi * (7)^2 A = pi * 49 A = 49 pi square inches

OpenStudy (jewlzme17):

Hope this helps you

OpenStudy (anonymous):

yes it does very much

OpenStudy (jewlzme17):

well that's good glad i could help